OpenStudy (anonymous):

Solve: x − 9 = 13 −4 4 −22 22

OpenStudy (compassionate):

x = 22

OpenStudy (compassionate):

Remember to do the opposite: x − 9 = 13 x - 9 + 9 = 13 + 9 x = 22 ;)

OpenStudy (hba):

very difficult question

OpenStudy (compassionate):

Could possibly be -4. or even -22. Tough one!

OpenStudy (hba):

thanks for explaining @Compassionate You are a very intelligent guy

OpenStudy (anonymous):

Yeah It is difficult for those who do not know the addition property of equality..

OpenStudy (compassionate):

^ Haha, yea, I guess that's true.

OpenStudy (anonymous):

guess??/ But I am sure..